OpenStudy (anonymous):

help!?! The radius of a sphere and of a cylinder are the same. The diameter of the sphere and the height of the

OpenStudy (whpalmer4):

\[V_{sphere} = \frac{4}{3}\pi r^3\]\[V_{cone} = \frac{1}{3}\pi r^2 h\]We know in this case that the height of the cone, \(h\) is the radius \(r\) so that becomes \[V_{cone} = \frac{1}{3}\pi r^3\] With that in mind, what's your answer?

OpenStudy (whpalmer4):

Oh, and the volume of the cylinder: \[V_{cylinder} = \pi r^2h = 2\pi r^3\]because the height \(h = 2r\)

OpenStudy (anonymous):

B?

OpenStudy (whpalmer4):

\[\frac{4}{3}\pi r^3 = 2\pi r^3 - 2(\frac{1}{3}\pi r^3)\]We can see that \(\pi r^3\) is a common factor to all of that, leaving us with \[\frac{4}{3} = 2 - \frac{2}{3}\]if we convert the 2 to 6/3, we see that the equation is correct, so answer B is correct. Good job!

OpenStudy (anonymous):

Thank you :)

OpenStudy (whpalmer4):

Hopefully you know those formulas well enough to recall them from memory. I'll confess I had to look up the formula for the cone :-)

OpenStudy (anonymous):

hopefully! lol